Three current carrying wires of length 5m are oriented perpendicularto the page and separated by .2m each. What is the net magnetic field at wire 2 from wires 1&3. What is the net force on wire 2 resulting from magnetic fields wires 1&3


perpendicular to the page and separated by from wi
Answer #1


Field due to currnt carrying wire at a distance d is B = mu_o*I/(2*pi*d)


so due to wire 1,magnetic field is B1 = mu_o*I1/(2*pi*d1) = (2*10^-7*2)/(2) = 2*10^-7 T downwards

due to wire 2 ,B2 = mu_o*I1/(2*pi*d2) = (2*10^-7*6)/2 = 6*10^-7 T upwards

net magnetic field is B = B2-B1 = (6-2)*10^-7 = 4*10^-7 T upwards

Net force on wire 2 is F = i*L*B*sin(90) = 4*5*4*10^-7*sin(90) = 8*10^-6 N